Consider a process consisting of three resources. Assume there exists unlimited demand for the product, and that all activities are always performed in the following sequence. 

• Resource 1 has a processing time of 7 minutes per unit. 

• Resource 2 has a processing time of 4 minutes per unit. 

• Resource 3 has a processing time of 6 minutes per unit. 

All three resources are staffed by one worker and each worker gets paid $10 per hour. 


a. What is the cost of direct labor? 

b. What is the labor content? 

c. What is the average labor utilization? percent

Do not round your intermediate calculations.

d. Assume the demand rate is 18 units per hour. What is the takt time? 

e. Assume the demand rate is 18 units per hour. What is the target manpower? 

f. If one additional worker could be hired, to which resource should the additional resources be assigned? Calculate the process capacity. 

g. If they could take 1 minute of task time from one resource and give it to another resource (so the labor content remains the same), What would be the new capacity of the process (units per hour)? Assume that there is still only 1 worker per station. 

h. What would be the capacity of the process if they integrated the work so that all workers do all three tasks (units per hour)? Assume that there is still only 1 worker per station. 
Answer #1

Output of resource 1 = 60/7 =8.5714 per hour

Output of resource 2 = 60/4 =15 per hour

Output of resource 3 = 60/6 =10 per hour

Resource 1 is bottleneck which restricts the output to 8.5714 per hour

(a) Cost of direct labour = 3x10 =30 per hour

  Cost of direct labour per unit = 30/8.5714 =3.5

(b) Labour content = 7+4+6 =17 min

(c) Utilisation of 1 = 8.5714/8.5714 =1

Utilisation of 2 = 8.5714/ 15 =0.5714

Utilisation of 3 = 8.5714/10 =0.8571

Avg utilisation = 0.8095

(d) Takt time = time available / units to be made = 60/18 =3.33 minutes

(e) For making 18 minutes per hour.

Resource 1 needed = 18/8.5714 = 3

Resource 2 needed = 18/15 =2
Resource 3 needed = 18/10 =2

Total manpower = 3+2+2=7

Note - As per policies, I can answer first 4 parts of a question. Inconvenience is regretted.